# Algebraically Skew Embedding of Curves

### v1

@@ 2.2 The basic upper bound of $\mathsf{msdim}\,X$
for any smooth points $x,y\in X$ and any smooth point $z\in\left\langle x,y\right\rangle\subseteq\mathsf{Sec}\,X$. This implies
@@
2.3 The degeneracy filtration of $\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N)$ and the lower bound of $\mathsf{msdim}\,X$
In this section we introduce and study the degeneracy filtration $D_{r}(n,N)$’s of $\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N)$, which is a series of varieties that help to measure how much a projective subvariety fails to be skew. In particular, we obtain that $3n\leq\mathsf{msdim}\,X$ in Corollary 2.5, which is a lower bound stronger than the real smooth case.
@@
We also write $\mathbb{D}_{r-1}=\mathbb{D}_{r}(n-1,N-1)=\mathbb{D}_{r-1}(\mathbb{P}(V))$ for the subvariety in ${\mathbb{G}\mathrm{r}}(n-1,\mathbb{P}(V))\times{\mathbb{G}\mathrm{r}}(n-1,\mathbb{P}(V))$ corresponding to $D_{r}(n,N)$.
@@
Proof.
Let $X\subseteq\mathbb{P}^{N}$. Let $\gamma:X\to{\mathbb{G}\mathrm{r}}(n,N)$ be the Gauss map of $X$ and set $\Gamma=\Gamma(X):=\gamma(X)\times\gamma(X)\subseteq{\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$.
Since, as subvarieties in ${\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$, we have $\mathsf{codim}\,\Gamma=2(n+1)(N-n)-2n$ and $\mathsf{codim}\,\mathbb{D}_{1}(n,\mathbb{P}^{N})=N-2n$, the expected dimension $d$ of the intersection $\Gamma\cap\mathbb{D}_{1}(n,\mathbb{P}^{N})$ is
$d=\mathsf{dim}\,({\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N))-\mathsf{codim}\,\Gamma-\mathsf{codim}\,\mathbb{D}_{1}(n,\mathbb{P}^{N})=4n-N.$
Observe that $X$ is skew in $\mathbb{P}^{N}$ if and only if, set theoretically, $\mathbb{D}_{1}(n,\mathbb{P}^{N})\cap\Gamma=\Delta_{\Gamma}$, where $\Delta_{\Gamma}$ is the diagonal of $\Gamma=\gamma(X)\times\gamma(X)$. Since $\mathsf{dim}\,\Delta_{\Gamma}=n$, we must have $n\geq d=4n-N$ for $X\subseteq\mathbb{P}^{N}$ to be skew. Thus $\mathsf{msdim}\,X\geq 3n$. ∎
Now, we compute the Schubert class of $D_{1}$ in $\mathsf{A}(\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N))$. Recall that by the standard Schubert calculus the Chow ring $\mathsf{A}(\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N))$ has basis $\sigma_{a}\otimes\sigma_{b}$’s, where $\sigma_{a}$’s and $\sigma_{b}$’s are Schubert cycles in $\mathsf{A}(\mathsf{Gr}(n,N))$; $a=(a_{1},\cdots,a_{n})$ and $b=(b_{1},\cdots,b_{n})$ are decreasing sequences of integers between $N-n$ and 00.
@@
By [Ful98, Theorem 14.4], we then obtain $[D_{1}]=c_{N-2n+1}(\mathcal{V}-(\mathcal{S}_{1}\oplus\mathcal{S}_{2}))$. Recall that $\mathcal{Q}_{i}\oplus\mathcal{S}_{i}=\mathcal{V}$ and $c(\mathcal{Q})=1+\sigma_{1}+\sigma_{2}+\cdots+\sigma_{N-n}$. Thus
@@
3 The blowup of the Grassmannian product and the skew curve in $\mathbb{P}^{3}$
From the discussion in the previous section, we know that a variety $X\subseteq\mathbb{P}^{N}$ is skew if and only if $(\Gamma\cap\mathbb{D}_{1})\setminus\Delta_{G}=\varnothing$, where $\Gamma$ and $\mathbb{D}_{1}$ are defined as in Corollary 2.5 and $\Delta_{G}$ is the diagonal of ${\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$. To handle the intersection of $\Gamma$ and $D_{1}$ along $\Delta_{G}$, we consider the proper transforms of $\mathbb{D}_{1}$ and $\Gamma$ in the blowup $B:=\mathsf{Bl}_{\Delta_{G}}{\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$.
In this section, we first study the structure of the Chow ring $\mathsf{A}(B)$. Then we compute the proper transforms of $D_{1}$ and $\Gamma$ for the case of $X$ being a curve in $\mathbb{P}^{N}$ and apply these tools to classify the skew curve in $\mathbb{P}^{3}$.
3.1 The structure of $\mathsf{A}(B)$
@@ 3.2 The class of $\widetilde{D}_{1}$
$[\widetilde{D}_{1}(2,4)][E]=j_{*}(2\overline{\sigma}_{1}-n_{0}\zeta)$ (7)
@@
$[\widetilde{D}_{1}(2,5)][E]=j_{*}(3\overline{\sigma}_{2}+\sigma_{11}-n_{0}\zeta^{2}-n_{1}\overline{\sigma}_{1}\zeta).$ (8)
@@
$\widetilde{D}_{1}\cap E=\left\{(x,[v]):\mathsf{rank}\,v<n\right\}=\left\{(x,[v]):\mathsf{rank}\,\phi_{(x,[v])}\leq n-1\right\},$ (9)
@@
$\left[\widetilde{D}_{1}(2,4)\cap E\right]=$ $j_{*}\left(\left\{c(\pi_{E}^{*}\mathcal{Q})s(\mathcal{O}_{E}(-1)\otimes\pi_{E}^{*}\mathcal{S})\right\}_{1}\right)$ (10)
@@
$\left[\widetilde{D}_{1}(2,5)\cap E\right]=$ $j_{*}\left(\left\{c(\pi_{E}^{*}\mathcal{Q})s(\mathcal{O}_{E}(-1)\otimes\pi_{E}^{*}\mathcal{S})\right\}_{2}\right)$ (11)
@@
3.3 The class of $\widetilde{\Gamma}$
We now turn our attention to the class of $\Gamma=\Gamma(X):=\gamma(X)\times\gamma(X)$ and its proper transform $\widetilde{\Gamma}=\widetilde{\Gamma}(X):=\pi^{-1}_{*}\Gamma(X)$ in $\mathsf{A}(B)$, where $X\subseteq\mathbb{P}^{N}$ is an $n$-dimensional projective variety and $G:=\mathsf{Gr}(n+1,N+1)={\mathbb{G}\mathrm{r}}(n,N)$. In particular, we focus on the case of $X$ being a smooth curve.
To start, we know that $[\widetilde{\Gamma}]=\pi^{*}([\Gamma])+j_{*}(\nu)$ for some $\nu\in\mathsf{A}_{\mathsf{dim}\,\Gamma}(E)$. Recall that $[\Gamma]=[\gamma(X)]\otimes[\gamma(X)]$. When $X$ is a nondegenerate curve, we have $[\gamma(X)]=d^{\vee}\sigma_{N-1,N-2}$ for some integer $d^{\vee}$. Since the Gauss image $\gamma(X)$ of a nondegenerate curve $X$ is again a curve, we have $\nu\in\mathsf{A}_{\mathsf{dim}\,\Gamma}(E)=\mathsf{A}_{2}(E)=\mathsf{A}^{4N-7}(E)$. Furthermore, by Lemma 3.3 we see that $\mathsf{A}(E)$ consists of the polynomials in $\pi_{E}^{*}(\mathsf{A}(\Delta_{G}))[\zeta]$ with degrees $\leq 2N-2$. Thus we have
@@ 3.3 The class of $\widetilde{\Gamma}$
$[E][\widetilde{\Gamma}]=$ $\left((d^{\vee})^{2}\pi^{*}\left(\sigma_{N-1,N-2}\otimes\sigma_{N-1,N-2}\right)+\sum_{2N-5\leq|a|\leq 4N-7}m_{a}j_{*}\left(\overline{\sigma}_{a}\zeta^{4N-7-|a|}\right)\right)j_{*}(1).$ (12)
@@
Given any $n$-dimensional smooth projective variety $X\subseteq\mathbb{P}^{N}$ with its Gauss map $\gamma:X\to{\mathbb{G}\mathrm{r}}(n,N)$ a diffeomorphism to its image. Let $E,\Gamma$, and $\widetilde{\Gamma}$ be defined as above. Then, under the identification $E\cong\mathbb{P}(\mathcal{T}_{G})$ from Lemma 3.4, we have $\widetilde{\Gamma}\cap E\cong\mathbb{P}(d\gamma(\mathcal{T}_{X}))=\mathbb{P}(\mathcal{T}_{\gamma(X)})$, where $d\gamma$ is the total derivative of $\gamma$.
@@
Let $p:\mathcal{T}_{G\times G}|_{\Delta_{G}}\to\mathcal{N}_{\Delta_{G},G\times G}:=\mathcal{T}_{G\times G}|_{\Delta_{G}}/\mathcal{T}_{\Delta_{G}}$ be the quotient map. Observe that, by the construction of blowup, $\widetilde{\Gamma}\cap E$ is the projectivization of $p(\mathcal{T}_{\Gamma}|_{\Delta_{G}\cap\Gamma})=p(\mathcal{T}_{\Gamma}|_{\Delta_{\Gamma}})$, where $\Delta_{\Gamma}:=\Gamma\cap\Delta_{G}$ is the diagonal of $\Gamma:=\gamma(X)\times\gamma(X)$. That is, $\widetilde{\Gamma}\cap E$ is the projectivization of the vector bundle
$p(\mathcal{T}_{\Gamma}|_{\Delta_{\Gamma}})=\left\{((x,x),p(0,v)):x\in\gamma(X),v\in\mathcal{T}_{x,\gamma(X)}\right\}$
@@
Let $\iota:G\to\Delta_{G}\subseteq G\times G$ be the isomorphism from $G$ to the diagonal $\Delta_{G}$ of $G\times G$, and $f:\iota^{*}\mathcal{N}_{\Delta_{G},G\times G}\to\mathcal{T}_{G}$ be the isomorphism of bundles defined in Lemma 3.4. Then we have $p(\mathcal{T}_{\Gamma}|_{\Delta_{\Gamma}})=f^{-1}(\mathcal{T}_{\gamma(X)})$, where $\mathcal{T}_{\gamma(X)}=d\gamma(\mathcal{T}_{X})$ is treated as a subbundle of $\mathcal{T}_{G}|_{\gamma(X)}$. Thus $\widetilde{\Gamma}\cap E\cong\mathbb{P}(d\gamma(\mathcal{T}_{X}))$. ∎
We proceed with the identifications $E\cong\mathbb{P}(\mathcal{T}_{G})$ and $\widetilde{\Gamma}\cap E\cong\mathbb{P}(\mathcal{T}_{\gamma(X)})$.
@@
where $\pi_{X}$ is the projection map and $\widetilde{\gamma}:\mathbb{P}(\gamma^{*}\mathcal{T}_{G})\to E$ is the map induced by the natural embedding from the pullback bundle $\gamma^{*}\mathcal{T}_{G}$ to $\mathcal{T}_{G}$.
@@
$[E\cap\widetilde{\Gamma}]=$ $j_{*}\left([\gamma(X)]\zeta^{2N-3}+\pi_{E}^{*}(c_{1}(\mathcal{T}_{G})[\gamma(X)]+\gamma_{*}(K))\zeta^{2N-4}\right)$ (18)
$=$ $d^{\vee}j_{*}\left(\overline{\sigma}_{N-1,N-2}\zeta^{2N-3}\right)+j_{*}\left(\left(d^{\vee}\overline{\sigma}_{N-1,N-2}\pi_{E}^{*}(c_{1}(\mathcal{T}_{G}))+(2g-2)\overline{\sigma}_{N-1,N-1}\right)\zeta^{2N-4}\right).$

### v4

@@ 2.2. The basic upper bound of $\mathsf{msdim}\,X$
for any generic smooth points $x,y\in X$ and any smooth point $z\in\left\langle x,y\right\rangle\subseteq\mathsf{Sec}\,X$. This implies
@@
Note that the upper bound above can be deduced from [BC21] as well.
2.3. The degeneracy filtration of $\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N)$ and the lower bound of $\mathsf{msdim}\,X$
In this section, we introduce and study the degeneracy filtration $D_{r}(n,N)$’s of $\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N)$, which is a series of varieties that help to measure how much a projective subvariety fails to be skew. In particular, we obtain that $3n\leq\mathsf{msdim}\,X$ in Corollary 2.6, which is a lower bound stronger than the real smooth case.
@@
We also write $\mathbb{D}_{r-1}=\mathbb{D}_{r-1}(n-1,N-1)=\mathbb{D}_{r-1}(\mathbb{P}(V))$ for the subvariety in ${\mathbb{G}\mathrm{r}}(n-1,\mathbb{P}(V))\times{\mathbb{G}\mathrm{r}}(n-1,\mathbb{P}(V))$ corresponding to $D_{r}(n,N)$.
@@
Let $C\subseteq{\mathbb{G}\mathrm{r}}(n,N)$ be an $m$-parameter family of $n$-dimensional linear subspaces in $\mathbb{P}^{N}$. Suppose that $C$ is skew, that is, $P_{1}\cap P_{2}=\varnothing$ for any $[P_{1}]\neq[P_{2}]\in C$, then $2n+m\leq N$.
Proof.
Set $\Gamma:=C\times C\subseteq G\times G:={\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$.
Since, as subvarieties in $G\times G$, we have $\mathsf{codim}\,\Gamma=2(n+1)(N-n)-2m$ and $\mathsf{codim}\,\mathbb{D}_{0}(n,\mathbb{P}^{N})=N-2n$, the expected dimension of the intersection $\Gamma\cap\mathbb{D}_{0}(n,\mathbb{P}^{N})$ is
$\mathsf{dim}\,({\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N))-\mathsf{codim}\,\Gamma-\mathsf{codim}\,\mathbb{D}_{0}(n,\mathbb{P}^{N})=2n+2m-N.$
Observe that $C$ is skew if and only if, set theoretically, $\mathbb{D}_{0}(n,\mathbb{P}^{N})\cap\Gamma=\Delta_{\Gamma}$, where $\Delta_{\Gamma}$ is the diagonal of $\Gamma:=C\times C$. Since $\mathsf{dim}\,\Delta_{\Gamma}=m$, we must have $m\geq 2n+2m-N$ for $C$ to be skew. Thus $2n+m\leq N$. ∎
In particular, we have the following lower bound for $\mathsf{msdim}\,X$.
Corollary 2.6
@@
Now, we compute the Schubert class of $D_{1}$ in $\mathsf{A}(\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N))$. Recall that by the standard Schubert calculus the Chow ring $\mathsf{A}(\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N))$ has basis $\sigma_{a}\otimes\sigma_{b}$’s, where $\sigma_{a}$’s and $\sigma_{b}$’s are Schubert cycles in $\mathsf{A}(\mathsf{Gr}(n,N))$; $a=(a_{1},\cdots,a_{n})$ and $b=(b_{1},\cdots,b_{n})$ are partitions in decreasing order that consist of integers between $N-n$ and 00.
@@
By [Ful98, Theorem 14.4], we then obtain $[D_{1}]=c_{N-2n+1}(\mathcal{V}-(\mathcal{S}_{1}\oplus\mathcal{S}_{2}))$. Recall the sequence $0\to\mathcal{S}_{i}\to\mathcal{V}\to\mathcal{Q}_{i}\to 0$ and $c(\mathcal{Q})=1+\sigma_{1}+\sigma_{2}+\cdots+\sigma_{N-n}$. Thus
@@
3. The blowup of the Grassmannian product and the skew curve in $\mathbb{P}^{3}$
From the discussion in the previous section, we know that a variety $X\subseteq\mathbb{P}^{N}$ is skew if and only if $(\Gamma\cap\mathbb{D}_{0})\setminus\Delta_{G}=\varnothing$, where $\Gamma$ and $\mathbb{D}_{0}$ are defined as in Corollary 2.6 and $\Delta_{G}$ is the diagonal of ${\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$. To handle the intersection of $\Gamma$ and $D_{1}$ along $\Delta_{G}$, we consider the proper transforms of $\mathbb{D}_{0}$ and $\Gamma$ in the blowup $B:=\mathsf{Bl}_{\Delta_{G}}{\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$.
In this section, we study the structure of the Chow ring $\mathsf{A}(B)$ and compute the proper transforms of $D_{1}$ and $\Gamma$ for the case of one-parameter family of lines in $\mathbb{P}^{N}$. Then we apply these tools to study the skewness of one-parameter family of lines and the skew curve in $\mathbb{P}^{3}$.
3.1. The structure of $\mathsf{A}(B)$
@@ 3.2. The class of $\widetilde{D}_{1}$
(7) $[\widetilde{D}_{1}(2,4)][E]=j_{*}(2\overline{\sigma}_{1}-n_{0}\zeta)$
@@
(8) $[\widetilde{D}_{1}(2,5)][E]=j_{*}(3\overline{\sigma}_{2}+\sigma_{11}-n_{0}\zeta^{2}-n_{1}\overline{\sigma}_{1}\zeta).$
@@
(9) $\widetilde{D}_{1}\cap E=\left\{(x,[v]):\mathsf{rank}\,v<n\right\}=\left\{(x,[v]):\mathsf{rank}\,\phi_{(x,[v])}\leq n-1\right\},$
@@
(10) $\left[\widetilde{D}_{1}(2,4)\cap E\right]=$ $j_{*}\left(\left\{c(\pi_{E}^{*}\mathcal{Q})s(\mathcal{O}_{E}(-1)\otimes\pi_{E}^{*}\mathcal{S})\right\}_{1}\right)$
@@
(11) $\left[\widetilde{D}_{1}(2,5)\cap E\right]=$ $j_{*}\left(\left\{c(\pi_{E}^{*}\mathcal{Q})s(\mathcal{O}_{E}(-1)\otimes\pi_{E}^{*}\mathcal{S})\right\}_{2}\right)$
@@
3.3. The class of $\widetilde{\Gamma}$
We now turn our attention to the class of $\Gamma=:=C\times C$ for some curve $C\subseteq G:=\mathsf{Gr}(n+1,N+1)={\mathbb{G}\mathrm{r}}(n,N)$ and its proper transform $\widetilde{\Gamma}=\widetilde{\Gamma}:=\pi^{-1}_{*}\Gamma$ in $\mathsf{A}(B)$. In particular, we focus on the case of $C=\gamma(X)$, the Gauss image of a smooth curve.
We start with the general case $\Gamma:=C\times C$, where $C\subseteq G$ is a smooth one-parameter family of lines in $\mathbb{P}^{N}$. We know that $[\widetilde{\Gamma}]=\pi^{*}([\Gamma])+j_{*}(\nu)$ for some $\nu\in\mathsf{A}_{\mathsf{dim}\,\Gamma}(E)$. Observe that $[\Gamma]=[C]\otimes[C]$, where $C=d^{\vee}\sigma_{N-1,N-2}$ for some integer $d^{\vee}$, and $\nu\in\mathsf{A}_{\mathsf{dim}\,\Gamma}(E)=\mathsf{A}_{2}(E)=\mathsf{A}^{4N-7}(E)$. Furthermore, by Lemma 3.3 we see that $\mathsf{A}(E)$ consists of the polynomials in $\pi_{E}^{*}(\mathsf{A}(\Delta_{G}))[\zeta]$ with degrees $\leq 2N-2$. Thus we have
@@ 3.3. The class of $\widetilde{\Gamma}$
(12) $[E][\widetilde{\Gamma}]=$ $\left((d^{\vee})^{2}\pi^{*}\left(\sigma_{N-1,N-2}\otimes\sigma_{N-1,N-2}\right)+\sum_{2N-5\leq|a|\leq 4N-7}m_{a}j_{*}\left(\overline{\sigma}_{a}\zeta^{4N-7-|a|}\right)\right)j_{*}(1).$
@@
Given any $n$-dimensional smooth family $C\subseteq G$ of lines in $\mathbb{P}^{N}$. Let $E$ and $\Gamma$ be defined as above, and $\widetilde{\Gamma}$ be the proper transform of $\widetilde{\Gamma}$. Then, under the identification $E\cong\mathbb{P}(\mathcal{T}_{G})$ from Lemma 3.4, we have $\widetilde{\Gamma}\cap E\cong\mathbb{P}(\mathcal{T}_{C})$.
@@
Let $p:\mathcal{T}_{G\times G}|_{\Delta_{G}}\to\mathcal{N}_{\Delta_{G},G\times G}:=\mathcal{T}_{G\times G}|_{\Delta_{G}}/\mathcal{T}_{\Delta_{G}}$ be the quotient map. Observe that, by the construction of blowup, $\widetilde{\Gamma}\cap E$ is the projectivization of $p(\mathcal{T}_{\Gamma}|_{\Delta_{G}\cap\Gamma})=p(\mathcal{T}_{\Gamma}|_{\Delta_{\Gamma}})$, where $\Delta_{\Gamma}:=\Gamma\cap\Delta_{G}$ is the diagonal of $\Gamma:=C\times C$. That is, $\widetilde{\Gamma}\cap E$ is the projectivization of the vector bundle
$p(\mathcal{T}_{\Gamma}|_{\Delta_{\Gamma}})=\left\{((x,x),p(0,v)):x\in C,v\in\mathcal{T}_{x,C}\right\}$
@@
Let $\iota:G\to\Delta_{G}\subseteq G\times G$ be the isomorphism from $G$ to the diagonal $\Delta_{G}$ of $G\times G$, and $f:\iota^{*}\mathcal{N}_{\Delta_{G},G\times G}\to\mathcal{T}_{G}$ be the isomorphism of bundles defined in Lemma 3.4. Then we have $p(\mathcal{T}_{\Gamma}|_{\Delta_{\Gamma}})=f^{-1}(\mathcal{T}_{C})$, where $\mathcal{T}_{C}$ is treated as a subbundle of $\mathcal{T}_{G}|_{C}$. Thus $\widetilde{\Gamma}\cap E\cong\mathbb{P}(\mathcal{T}_{C})$. ∎
We proceed with the identifications $E\cong\mathbb{P}(\mathcal{T}_{G})$ and $\widetilde{\Gamma}\cap E\cong\mathbb{P}(\mathcal{T}_{C})$.
@@
where $\gamma:C\hookrightarrow G$ is the embedding of $C$, $\pi_{C}$ is the projection map and $\widetilde{\gamma}:\mathbb{P}(\gamma^{*}\mathcal{T}_{G})\to E$ is the map induced by the natural embedding from the pullback bundle $\gamma^{*}\mathcal{T}_{G}$ to $\mathcal{T}_{G}$.
@@
(18) $[E\cap\widetilde{\Gamma}]=$ $j_{*}\left([C]\zeta^{2N-3}+\pi_{E}^{*}(c_{1}(\mathcal{T}_{G})[C]+\gamma_{*}(K))\zeta^{2N-4}\right)$
$=$ $d^{\vee}j_{*}\left(\overline{\sigma}_{N-1,N-2}\zeta^{2N-3}\right)+j_{*}\left(\left(d^{\vee}\overline{\sigma}_{N-1,N-2}\pi_{E}^{*}(c_{1}(\mathcal{T}_{G}))+(2g-2)\overline{\sigma}_{N-1,N-1}\right)\zeta^{2N-4}\right),$
where $K$ is the canonical class of the curve $C$.
Lemma 3.9
Let $C\subseteq G$ be a one-parameter family of lines in $\mathbb{P}^{N}$ of genus $g$. Then
  1. (i)
    for $N=3$,
    $[\widetilde{\Gamma}]=(d^{\vee})^{2}\pi^{*}(\sigma_{21}\otimes\sigma_{21})-j_{*}\left(d^{\vee}\overline{\sigma}_{21}\zeta^{2}+(4d^{\vee}+2g-2)\overline{\sigma}_{22}\zeta\right);$
  2. (ii)
    for $N=4$,
    $[\widetilde{\Gamma}]=(d^{\vee})^{2}\pi^{*}(\sigma_{32}\otimes\sigma_{32})-j_{*}\left(d^{\vee}\overline{\sigma}_{32}\zeta^{4}+(5d^{\vee}+2g-2)\overline{\sigma}_{33}\zeta^{3}\right),$
where $d^{\vee}$ is the coefficient of the class $[C]\in\mathsf{A}(G)$. In particular, if $X\subseteq\mathbb{P}^{N}$ be a degree $d$ genus $g$ smooth curve with its Gauss map $\gamma:X\to{\mathbb{G}\mathrm{r}}(1,N)$ a diffeomorphism to its image. where $d^{\vee}:=2d+2g-2$.
Proof.
(19) $[E\cap\widetilde{\Gamma}]=$ $d^{\vee}j_{*}\left(\overline{\sigma}_{21}\zeta^{3}\right)+j_{*}\left((d^{\vee}\overline{\sigma}_{21}(4\overline{\sigma}_{1})+(2g-2)\overline{\sigma}_{22})\zeta^{2}\right)$
(20) $[E\cap\widetilde{\Gamma}]=$ $(2d+2g-2)j_{*}\left(\overline{\sigma}_{32}\zeta^{5}\right)+j_{*}\left(\left((2d+2g-2)\overline{\sigma}_{32}(5\overline{\sigma}_{1})+(2g-2)\overline{\sigma}_{33}\right)\zeta^{4}\right)$
The lemma follows from comparing the coefficients above and the coefficients in (12) and the Lemma below. ∎
@@
3.4. Self-intersection of one-parameter families of lines in $\mathbb{P}^{3}$
Now that $[\widetilde{D}_{1}],[\widetilde{\Gamma}]$, and the blowup structure are established, we are ready to study the self-intersection of one-parameter families of lines in $\mathbb{P}^{3}$.